OpenStudy (itsbribro):

PLEASE HELP!!!! What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3?

OpenStudy (itsbribro):

@crazychickengirl123

OpenStudy (itsbribro):

@lmw

OpenStudy (anonymous):

The equation of the original line is y = 2x-4; this equation is in the form y = mx+c, where m is the slope of the line. As a result, the slope of the line is 2. The question requires the finding of the equation of the line that is perpendicular to the original line; as a result, we have to find the slope of the perpendicular line first. Since the slope of a line * the slope of its perpendicular line = -1; therefore, 2 * the slope of the perpendicular line = -1 the slope of the perpendicular line = -1/2. Given the perpendicular line passes through the point (-6,2), by using point slope form, which is y-y1 = m(x-x1) By substituting the point (-6,2) into the point slope form, y-2 = (-1/2)*[x-(-6)] y-2 = (-1/2)*(x+6) y-2 = (-1/2)(x)-3 y = (-1/2)(x)-1; Therefore, the slope-intercept form of the perpendicular line is y = (-1/2)(x)-1. source: http://answers.yahoo.com/question/index?qid=20081114094710AAQ5XEz

OpenStudy (anonymous):

this help??

OpenStudy (itsbribro):

kinda umm thx lol